# Invariant Federated Learning for Edge Intelligence: Mitigating Heterogeneity and Asynchrony via Exit Strategy and Invariant Penalty

### v2

@@ Abstract
This paper provides an invariant federated learning system for resource-constrained edge intelligence. This framework can avoid the impact of heterogeneity and asynchrony by exit strategy and invariant penalty. We decompose local information into two orthogonal components to measure the contribution or impact of heterogeneous and asynchronous clients. We propose that the exit of abnormal clients can guarantee the effect of the model on most clients. Meanwhile, to ensure the models’ performance on exited abnormal clients and those who lack training resources, we propose Federated Learning with Invariant Penalty for Generalization (FedIPG) based on the invariant orthogonal decomposition of parameters. Theoretical proof shows that FedIPG reduces the Out-Of-Distribution prediction loss without increasing the communication burden. The performance of FedIPG combined with an exit strategy is tested empirically in multiple scales using four datasets. It shows our system can enhance In-Distribution performance and outperform the state-of-the-art algorithm in Out-Of-Distribution generalization while maintaining model convergence. Additionally, the results of the visual experiment prove that FedIPG contains preliminary causality in terms of ignoring confounding features.
@@
Invariant learning, Asynchronous Federated Learning, Heterogeneous environment, Out-of-Distribution Generalization, Edge intelligence
@@ I Introduction
As smart devices become increasingly lightweight and the demand for terminal task computing grows, the intelligence of edge devices has emerged as a prominent trend. With large volumes of data being collected and stored locally on these devices, there is an escalating need for a distributed learning framework to address communication constraints while ensuring robust privacy protection. In response to these challenges, federated learning (FL) has emerged as a key area of research.
@@
Federated learning is a distributed computing paradigm designed to preserve privacy while enabling collaborative learning. The central tenet of FL is that data remains on local devices, with minimal interaction with a central server or other clients[1, 2]. Previous research has shown that, under ideal conditions, FL can achieve performance comparable to centralized learning. However, these ideal conditions are often challenging to attain in edge intelligence contexts, as shown in Fig.1. First, data collected from diverse sources typically exhibit significant heterogeneity and environmental interference. Second, in sophisticated scenarios involving edge intelligence, there exists a situation where communication and computational capabilities are unevenly distributed, ensuring synchronization within edge intelligence systems poses a considerable challenge. Both data heterogeneity and asynchrony can substantially degrade the performance of FL in such environments[3, 4].
While much of the existing research has focused on decreasing the harm heterogeneous and asynchronous clients caused in federated aggregation, inspired by invariant learning, we introduce a novel perspective that suggests heterogeneous and asynchronous clients exit the training process. Specifically, the model should be only employed on those aberrant clients after training. Our previous work has demonstrated that when latency is high, it is often more advantageous to allow computationally inefficient clients with poor communication environments to withdraw from the federated training process[5]. In this article, we decompose the local optimal parameters of clients into two orthogonal parameter vectors: the globally practical invariant component and the locally effective environmental heterogeneous component[6]. We proved that, as the size of the heterogeneous component grows and the delay increases, the anomalous clients are more likely to provide a negative effect rather than a positive contribution, which means the exit strategy can protect the model from disturbance by heterogeneous and asynchronous clients. Suppose models’ generalization capability can guarantee the performance on aberrant clients’ prediction. In that case, the withdrawal of aberrant clients from global aggregation can enhance the model’s performance on regular clients, which account for the majority, without harming the rights of the few exited clients. Although the withdrawal of anomalous clients indirectly addresses the harm caused by heterogeneity and asynchrony, leading to better-performing models for In-Distribution (ID) clients, the performance of Out-Of-Distribution (OOD) clients that exit the training process is no longer guaranteed. Additionally, clients lacking essential resources for training, such as datasets or computational power, still require the model to have inference capabilities on out-of-distribution datasets. In such cases, clients may bypass training altogether and directly utilize the trained model, necessitating the model to exhibit strong OOD generalization capabilities. The issue of OOD generalization has also been extensively studied in centralized learning [7, 8, 9]. One notable approach garnered significant attention is Invariant Learning[10]. The core lies in decoupling data features into invariant and environmental features. Invariant features are considered stable and effective across diverse conditions, while environmental features, such as the background of an image, may introduce noise or instability[11]. For instance, in wildfire detection, flames and temperature represent invariant features. At the same time, the background (e.g., smoke or fog) serves as an environmental feature that may interfere with classification[12]. Several variants of centralized schemes have been proposed to leverage this relationship. However, there has been limited research on applying a centralized invariant scheme within the context of FL systems[13, 14].
@@
Inspired by Invariant Risk Minimization[15], We proposed FedIPG to realize OOD generalization in edge intelligence scenarios to satisfy the prediction needs of exited clients in the exit strategy we proposed. FedIPG employs gradients to approximate environmental heterogeneous components and constructs an invariance penalty term. This penalty term reduces the impact of environment-specific parameters, facilitating learning globally invariant parameters. This model paradigm complements the potential distribution shift and fairness violations caused by the exit strategy, ensuring that clients not participating in training have basic performance guarantees. By combining the exit strategy with FedIPG, FL can allocate more communication resources to clients with strong computational capabilities in edge learning scenarios characterized by uneven computational resource distribution and limited communication resources, thereby enabling the training of a globally invariant model. The key contributions and innovations of this paper are as follows:
@@
  • We introduce the invariant learning paradigm from centralized learning into the federated learning framework and propose a federated learning model, Federated Learning with Invariant Penalty for Generalization, which enhances OOD generalization.
  • We empirically prove that FedIPG with exit strategy improves the model’s performance in both ID data and the OOD affine space without impeding the convergence of the model. Additionally, it is found that FedIPG has the preliminary ability to resist confounding in causal learning.
@@ II Related Work
In the field of federated learning, numerous studies have addressed non-IID problems [13, 16] and asynchronous federated learning [17, 18]. Several works have focused on selecting a specific subset of clients to mitigate the impact of outdated gradients [19, 20]. However, these approaches often necessitate real-time monitoring of clients’ channel conditions, which are highly volatile in edge-intelligent environments [21]. Furthermore, the delays resulting from computational constraints cannot be effectively resolved solely through client selection strategies. To date, no research has theoretically quantified the potential harm caused by the participation of heterogeneous and asynchronous clients in training by parameter orthogonal decomposition.
In centralized learning, supervised model learning for OOD generalization normally needs environment labels to improve the learning of features or representations. Invariant Causal Prediction (ICP)[22] is an important approach towards inferring causal structure via invariance property[23]. However, in most scenarios, the assumption that ICP is based on is violated, so the effectiveness of ICP depends highly on the quality of availability of environment[24]. The Causal Semantic Generative model(CSG) identifies the semantic factor to guarantee the boundedness of OOD generalization error[25] in the image classification task. Causal Invariant Transformation(CIT) [26] uses transformations to modify non-causal features while leaving the causal part unchanged. Risk Extrapolation(REx) and its variant, MM-REx, trade-off system’s robustness to causally induced distribution shifting and covariate to reinforce the stability to OOD input[27]. Another solution to OOD challenges is to proficiently extract invariant features from data and establish accurate mappings between these features and corresponding labels. IRM algorithm is a novel learning paradigm that estimates nonlinear, invariant, causal predictors from multiple training environments that can effectively fulfill the above requirements. However, IRM is a highly non-convex bi-level objective, [15] proposed a practical vision to reduce the optimization difficulty. Although the functionality of IRM seems promising[28], compared to the result from ERM, its function will fail in some specific scenarios[29]. However, most of these solutions to OOD generalization, including IRM, are discussed based on the centralized learning structure and do not consider the limitations caused by data isolation.
Some current studies focus on domain generalization problems in FL, such as FedADG[30] and FedSR[31]. These methods are based on the data distribution and require additional models or calculations to align the distribution. Distributed masking function and invariant learning are introduced into FL in FedGen[32], which outperforms current FL in efficiency, accuracy, and generalization in various contexts. However, the FedGen model focuses on sequential data, which differs from our approach. FedIIR[33] is the latest research on invariant FL. From the perspective of implicit, FedIIR learns invariant relationships for OOD generalization by exchanging the gradient of the classifier at the beginning of each round of communication. FedIIR performs excellently in federated systems, but it has additional communication requirements, which are not applicable in the scenario of edge intelligence.
@@ III-A The Federated System Model with Exit Strategy
Algorithm 1 Federated Learning with Invariant Penalty for Generalization
1:Initialization:
2:Set $t=1$
3:The center sends the initialized model parameters $w^{1}$ to all nodes
4:Center Side:
5:for $t<T+1$ do
6:     Aggregation: $w^{t+1}=\sum_{i\in\mathcal{I}_{ID}}\frac{n_{i}}{\sum_{j\in\mathcal{I}_{ID}}n_{j}}w^{t}_{i}$.
7:     Distribute $w^{t+1}$ to the client
8:     Set $t=t+1$
9:end for
10:Client Side:
11:for $t<T+1$ do
12:     for all clients: $i={1,2,..,N}$ do
13:         Receive $w^{t}$
14:         Compute $w^{t}_{i}=w^{t}-\eta\nabla f_{i}(w^{t})$
15:         Upload $w^{t}_{i}$ to the server
16:     end for
@@
18:for $t=T+1$ do
19:     for all clients: $i\in\mathcal{I}_{all}$ do
20:         Apply $w^{t}_{i}$ to prediction tasks
21:     end for
@@
III-B The Profile of the Invariant Schemes and Notations
@@
Our objective is to steer the parameters towards emphasizing invariant features that maintain a causal relationship with the task across diverse client environments, thereby facilitating generalization to OOD data, as shown in figure 2. To achieve OOD generalization, the model needs to decouple the data $X$ into two independent features: the invariant features $X_{I}$ and the heterogeneous features $V$. Inspired by previous researchers [36] , we decompose the optimal parameter vector of each client into two orthogonal vectors, $w^{I}$ and $w^{e}_{i}$. Here, $w^{I}$ corresponds to $X_{I}$, which remains consistent across all clients, while $w^{e}_{i}$ corresponds to $V$, which varies from client to client. The orthogonality between $w^{I}$ and $w^{e}_{i}$ reflects the independence between $X_{I}$ and $V$
$w^{I}=w^{*}_{i}-w^{e}_{i},\ w^{e}_{i}\perp w^{I}$ (6)
Consequently, our objective shifts from identifying $X_{I}$ to training a parameterized representation $w^{I}$ of the invariant features within the federated system. Since $-\nabla R_{i}(w)$ can be interpreted as pointing toward the optimal parameter when the current parameter is $w$, i.e., $-\nabla R_{i}(w)\approx w^{*}_{i}-w$. If $w=w^{I}$, then $-\nabla R_{i}(w^{I})\perp w^{I}$. To identify the global invariant parameter, we need to ensure that the following equality holds
@@
Based on the above reasoning, the loss function of client $i$ in the FedIPM system is defined as:
@@
Algorithm 1 shows the pseudo-code of FedIPM. In conclusion, each client locally takes a step of gradient descent based on the current global parameter and the invariant loss function calculated with local data. The server then takes a weighted average of the local parameter as the global parameter.
@@
IV The basis of Aberrant Clients’ Exit Strategy
@@ IV-A Theoretical Measure for Clients’ Contribution
In this section, we define the contribution of a client’s participation in a round of aggregation. When there is a probability that a client’s contribution is negative, it implies that the client’s participation may have a detrimental effect on the model’s performance. By orthogonally decomposing parameters, the contributions of heterogeneous and asynchronous clients are quantified.
@@ IV-B Contribution Analysis for Aberrant clients
As a result, the probability of an aberrant client’s contribution being negative grows with the size of the heterogeneous component and the delay, suggesting that excluding such clients benefits aggregation. Prohibiting aberrant clients from participating can effectively mitigate this negetive contribution.
@@ V-B Convergence Analysis of Invariant Asynchronous Federated Learning Algorithm
We next formally prove that the model with invariant penalty results in an aggregated model by the above Assumptions and Lemmas.
@@
Based on the above Lemma, the convergence of loss function FedIPM without isolation of data can be proved in Appendix-B Based on the above conclusions, we have established that the loss function incorporating the penalty term is convex and smooth. Additionally, in the centralized setting, grouping data by environment allows us to obtain the optimal invariant parameters under the current penalty term using gradient descent. Next, we will shift our focus to the context of edge intelligence, analyzing the convergence properties of invariant learning within a federated framework.

### v3

@@ Abstract
This paper provides an invariant federated learning system for resource-constrained edge intelligence. This framework can mitigate the impact of heterogeneity and asynchrony via exit strategy and invariant penalty. We introduce parameter orthogonality into edge intelligence to measure the contribution or impact of heterogeneous and asynchronous clients. It is proved in this paper that the exit of abnormal edge clients can guarantee the effect of the model on most clients. Meanwhile, to ensure the models’ performance on exited abnormal clients and those who lack training resources, we propose Federated Learning with Invariant Penalty for Generalization (FedIPG) by constructing the approximate orthogonality of the invariant parameters and the heterogeneous parameters. Theoretical proof shows that FedIPG reduces the Out-Of-Distribution prediction loss without increasing the communication burden. The performance of FedIPG combined with an exit strategy is tested empirically in multiple scales using four datasets. It shows our system can enhance In-Distribution performance and outperform the state-of-the-art algorithm in Out-Of-Distribution generalization while maintaining model convergence. Additionally, the results of the visual experiment prove that FedIPG contains preliminary causality in terms of ignoring confounding features.
@@
Invariant learning, Asynchronous Federated Learning, Parameter Orthogonality, Heterogeneous environment, Out-of-Distribution Generalization, Edge intelligence
@@ I Introduction
As smart devices become increasingly lightweight and the demand for terminal task computing grows, the intelligence of edge devices has emerged as a prominent trend. With large volumes of data being collected and stored locally on these devices, there is an escalating need for a distributed learning framework to address communication constraints while ensuring robust privacy protection. In response to these challenges, federated learning (FL) has emerged as a key area of research.FL is a distributed computing paradigm designed to preserve privacy while enabling collaborative learning. The central tenet of FL is that data remains on local devices, with minimal interaction with a central server or other clients[1, 2].
@@
Although previous research has shown that, under ideal conditions, FL can achieve performance comparable to centralized learning, in edge intelligence contexts these ideal conditions are challenging to achieve, as shown in Fig.1. First, data collected from diverse sources typically exhibit significant heterogeneity and environmental interference. Second, in sophisticated scenarios involving edge intelligence, there exists a situation where communication and computational capabilities are unevenly distributed, ensuring synchronization within edge intelligence systems poses a considerable challenge. Both data heterogeneity and asynchrony can substantially degrade the performance of FL in such environments[3, 4]. While existing works have focused on mitigating the negative impacts of heterogeneous and asynchronous clients through improved aggregation schemes, our approach replenish a new perspective on modeling and analysis.
We propose an innovative framework where abnormal clients detected during the pretraining phase can autonomously withdraw from the training process, thereby alleviating the impacts of system heterogeneity and communication asynchrony. The trained model will be only employed on those aberrant clients after training. Our previous work has demonstrated that when latency is high, it is often more advantageous to allow computationally inefficient clients with poor communication environments to withdraw from the federated training process[5]. In this article, we decompose the local optimal parameters of clients into two parameter vectors: the globally practical invariant component and the locally effective environmental heterogeneous component[6]. The locally environmental nuisance parameters are required to be orthogonal to the parameter of invariant interest[7]. We theoretically prove that as the size of the heterogeneous component increases and delays become more pronounced, anomalous clients are likely to negatively impact global aggregation in FedAvg[8]. The proposed exit strategy thus protects the model from disturbances caused by such clients. Provided that the model can generalize well, their withdrawal can improve performance for the majority of regular clients without compromising the utility of the model for the few clients that exit.
Although the exit strategy enhances performance for In-Distribution (ID) clients, it raises concerns regarding Out-Of-Distribution (OOD) generalization for exited clients and clients who lacking sufficient data or computational resources but still requiring robust OOD inference. This challenge has been addressed in centralized learning  [9, 10, 11], with invariant learning emerging as a prominent solution [12]. Invariant learning decouples data features into invariant (stable and effective across conditions) and environmental components, where environmental features can introduce noise or instability [13]. For instance, in wildfire detection, flames and temperature represent invariant features, whereas the background (e.g., smoke or fog) can confound the classification [14]. Although several centralized schemes have leveraged this idea, applying it within FL remains limited [15, 16].
@@
Inspired by Invariant Risk Minimization[17] and parameter orthogonality mentioned earlier, we propose FedIPG to achieve OOD generalization in edge intelligence scenarios. FedIPG employs gradients to approximate environmental heterogeneous components and constructs an invariance penalty term to suppress the influence of environment-specific variations, thereby facilitating the learning of globally invariant parameters.This process, firmly grounded in parameter orthogonality, ensures that the environmental component does not undermine the estimation of invariant features, resulting in a robust model even for clients that exit the training process.By combining the exit strategy with FedIPG, our framework efficiently allocates communication resources to clients with stronger computational capabilities, leading to a globally invariant model that performs well on both ID and OOD data while safeguarding against fairness concerns. The key contributions and innovations of this paper are as follows:
@@
  • We introduce an invariant learning paradigm into the FL framework and propose a novel algorithm, Federated Learning with Invariant Penalty for Generalization (FedIPG), which enhances OOD generalization. Our approach is built upon the principle of parameter orthogonality, ensuring effective disentanglement of invariant and heterogeneous components.
  • We empirically demonstrate that FedIPG combined with the exit strategy improves performance on both ID and OOD data, outperforming baseline algorithms. Moreover, FedIPG shows promising resistance to confounding in causal learning scenarios.
@@ II Related Work
In the field of federated learning, numerous studies have addressed the challenges of non-IID data distributions [15] and asynchronous updates [18, 19]. To mitigate these issues, several works focus on selecting a specific subset of clients to reduce the negative effects of outdated gradients [20, 21]. However, such methods typically require real-time monitoring of clients’ channel conditions, a task that is particularly challenging in edge-intelligent environments, where the channel state is highly volatile [22]. Moreover, no prior work has provided a theoretical quantification of the potential harm caused by the participation of heterogeneous and asynchronous clients using a parameter orthogonal decomposition approach, a principle that is critical for ensuring the asymptotic independence of the estimates for the invariance parameter and the environmental heterogeneity parameter. This orthogonality minimizes interference between these two components, leading to more stable and robust statistical inferences [7].
In centralized learning, methods for out-of-distribution (OOD) generalization often rely on environment labels to guide feature or representation learning. For example, Invariant Causal Prediction (ICP) [23] leverages invariance properties to infer causal structures [13], though its effectiveness is highly dependent on the availability and quality of environment information [24]. Other approaches, such as the Causal Semantic Generative model (CSG) [25] and Causal Invariant Transformation (CIT) [26], respectively extract semantic factors to bound generalization errors and modify non-causal features while preserving causal ones. Methods like Risk Extrapolation (REx) and its variant, MM-REx, trade off robustness to distribution shifts with stability against covariate variations [27], while Invariant Risk Minimization (IRM) [17] proposes to estimate nonlinear invariant predictors from multiple environments. Despite the promise of IRM, its non-convex bi-level formulation poses significant optimization challenges, and its performance may degrade under certain scenarios compared to empirical risk minimization (ERM) [28, 29]. Importantly, these centralized methods do not account for the inherent data isolation present in FL systems.
In the federated domain, recent works such as FedADG [30] and FedSR [31] have focused on domain generalization through distribution alignment—usually at the cost of increased modeling complexity and additional computational overhead. Meanwhile, FedGen [32] employs distributed masking functions with invariant learning to achieve improvements in efficiency, accuracy, and generalization for sequential data, and FedIIR [33] implicitly learns invariant relationships through early-round gradient exchange. However, the latter incurs additional communication overhead, rendering it less suitable for edge-intelligent scenarios with constrained resources. In contrast to these works, our study uniquely quantifies the detrimental effects of including heterogeneous and asynchronous clients through a parameter orthogonal decomposition framework and investigates improving the generalization of the model on clients not involved in the training by using parameter orthogonality without any additional communication in formal training process.
@@ III-A The Federated System Model with Exit Strategy
Algorithm 1 Federated Learning with Invariant Penalty for Generalization with exit strategy
1:Pretraining: Identify $\mathcal{I}_{ID}\in\mathcal{I}_{all}$
2:Initialization:
3:Set $t=1$
4:The center sends the initialized model parameters $w^{1}$ to all nodes
5:Center Side:
6:for $t<T+1$ do
7:     Aggregation: $w^{t+1}=\sum_{i\in\mathcal{I}_{ID}}\frac{n_{i}}{\sum_{j\in\mathcal{I}_{ID}}n_{j}}w^{t}_{i}$.
8:     Distribute $w^{t+1}$ to the client
9:     Set $t=t+1$
10:end for
11:Client Side:
12:for $t<T+1$ do
13:     for all clients: $i={1,2,..,N}$ do
14:         Receive $w^{t}$
15:         Compute $w^{t}_{i}=w^{t}-\eta\nabla f_{i}(w^{t})$
16:         Upload $w^{t}_{i}$ to the server
@@
18:end for
19:for $t=T+1$ do
20:     for all clients: $i\in\mathcal{I}_{all}$ do
21:         Apply $w^{t}_{i}$ to prediction tasks
@@
III-B The Profile of the Parameter Orthogonality the Invariant Schemes
@@
Our objective is to steer the parameters towards emphasizing invariant features that maintain a causal relationship with the task across diverse client environments, thereby facilitating generalization to OOD data, as shown in figure 2. To achieve OOD generalization, the model needs to decouple the data $X$ into two independent features: the invariant features $X_{I}$ and the heterogeneous features $V$. Inspired by previous researchers [36], we decompose the optimal parameter vector of each client into two orthogonal vectors, $w^{I}$ and $w^{e}_{i}$. Here, $w^{I}$ corresponds to $X_{I}$, which remains consistent across all clients, while $w^{e}_{i}$ corresponds to $V$, which varies from client to client. The orthogonality between $w^{I}$ and $w^{e}_{i}$ reflects the independence between $X_{I}$ and $V$. According to parameters orthogonality theorem[7], we define $w^{I}$ and $w^{e}_{i}$ as
$E(\frac{\partial R_{i}(w)}{\partial w^{e}_{i}}\frac{\partial R_{i}(w)}{\partial w^{I}};w)=0,$ (6)
where $R_{i}(w^{I}+w^{e}_{i})=0$. Since $-\nabla R_{i}(w)$ can be interpreted as pointing toward the optimal parameter when the current parameter is $w$, i.e., $-\nabla R_{i}(w)\approx w^{*}_{i}-w$. So, we can rewrite the relationship between $w^{I}$ and $w^{e}_{i}$ as:
$\forall i:\ E(<-\nabla R_{i}(w^{I}),w^{I}>;w=w^{I}+w^{e}_{i})=0,$ (7)
If $w=w^{I}$, then $-\nabla R_{i}(w^{I})\perp w^{I}$. Consequently, our goal shifts to training a invariant parameter $w^{I}$ within the federated system. To identify the global invariant parameter, we need to ensure that the following equality holds
@@
Based on the above reasoning, the loss function of client $i$ in the FedIPG system is defined as:
@@
Algorithm 1 shows the pseudo-code of FedIPG. In conclusion, each client locally takes a step of gradient descent based on the current global parameter and the invariant loss function calculated with local data. The server then takes a weighted average of the local parameter as the global parameter.
@@
IV The basis of Aberrant Clients Exit Strategy
@@ IV-A Theoretical Measure for Clients’ Contribution
In this section, we define the contribution of a client’s participation in a round of aggregation in FedAvg. When there is a probability that a client’s contribution is negative, it implies that the client’s participation may have a detrimental effect on the model’s performance. By orthogonally decomposing parameters, the contributions of heterogeneous and asynchronous clients are quantified.
@@ IV-B Contribution Analysis for Aberrant clients
As the magnitude of a client’s heterogeneous component and its delay increase, the likelihood of its contribution being negative also rises. This observation motivates a scheme in which aberrant clients proactively opt out of the training process rather than waiting for the server to exclude them. Prior to formal training, we assume that all clients and the server jointly conduct a pre-training phase, modeled similarly to the design in [39], that assesses each client’s delay and data quality. Based on this self-assessment, clients can decide whether to participate in subsequent training rounds. Notably, similar contribution metrics can be incorporated into asynchronous or heterogeneous federated settings, indicating that our approach complements and extends existing research rather than contradicting it.
@@ V-B Convergence Analysis of Invariant Asynchronous Federated Learning Algorithm
We next formally prove that the model with invariant penalty results in an aggregated model by the above Assumptions.
@@
Based on the above conclusions, we have established that the loss function incorporating the penalty term is convex and smooth. Additionally, in the centralized setting, grouping data by environment allows us to obtain the optimal invariant parameters under the current penalty term using gradient descent which is proved in Appendix-B. Next, we will shift our focus to the context of edge intelligence, analyzing the convergence properties of invariant learning within a federated framework.